### v2

# Tail Distribution of Regret in Optimistic Reinforcement Learning

Abstract
We derive instance-dependent tail bounds for the regret of optimism-based reinforcement learning in finite-horizon tabular Markov decision processes with unknown transition dynamics. Focusing on a UCBVI-type algorithm, we characterize the tail distribution of the cumulative regret $R_{K}$ over $K$ episodes, rather than only its expectation or a single high-probability quantile. We analyze two natural exploration-bonus schedules: (i) a $K$-dependent scheme that explicitly incorporates the total number of episodes $K$, and (ii) a $K$-independent scheme that depends only on the current episode index. For both settings, we obtain an upper bound on $\mathbb{P}(R_{K}\geq x)$ that exhibits a distinctive two-regime structure: a sub-Gaussian tail starting from an instance-dependent scale $m_{K}^{\mathrm{mb}}$ up to a transition threshold, followed by a sub-Weibull tail beyond that point. We further derive corresponding instance-dependent bounds on the expected regret $\mathbb{E}[R_{K}]$. The proposed algorithm depends on a tuning parameter $\alpha$, which balances the expected regret and the range over which the regret exhibits a sub-Gaussian tail. To the best of our knowledge, our results provide one of the first comprehensive tail-regret guarantees for a standard optimistic algorithm in episodic reinforcement learning.
1 Introduction
We consider online Reinforcement Learning (RL) in finite-horizon Markov decision processes (MDPs) with unknown transition dynamics. In this setting, a learning agent interacts with an episodic environment for $K$ episodes of length $H$, and aims to minimize its cumulative regret with respect to an optimal policy. A long line of work has established sharp worst-case regret guarantees for model-based and model-free algorithms in tabular MDPs; see, for instance, the near-minimax optimality results for UCRL2 under the average-reward criterion (Jaksch et al., 2010) and for UCBVI and its variants in the episodic setting (Azar et al., 2017; Zanette and Brunskill, 2019). These results show that, up to logarithmic factors, the optimal worst-case regret scales as $\widetilde{\mathcal{O}}(\sqrt{HSAK})$, where $S$ and $A$ are the sizes of the state and action spaces.
More recently, there has been substantial progress on instance-dependent analysis in RL. Inspired by the classical gap-dependent theory of stochastic bandits, several works have derived regret bounds that depend on the suboptimality gaps of state–action pairs rather than only on $(S,A,H,K)$. For tabular MDPs, this includes non-asymptotic gap-dependent bounds for model-based algorithms (Simchowitz and Jamieson, 2019), fine-grained gap-dependent bounds via adaptive multi-step bootstrapping (Xu et al., 2021), and gap-dependent guarantees for model-free $Q$-learning and its refinements (Yang et al., 2021; Velegkas et al., 2022; Zheng et al., 2024). These results provide a more refined characterization of the regret in MDPs by incorporating an instance-dependent $Q^{*}$-gap, which measures the separation between optimal and suboptimal actions at each state.
At the same time, the tail behavior of regret in RL is still poorly understood. Most existing analyses deliver bounds of the form
$\mathbb{P}\!\big(R_{K}\geq f(K,\delta)\big)\;\leq\;\delta,$
for a fixed confidence level $\delta\in(0,1)$, typically by combining a union bound with martingale concentration inequalities. While such high-probability guarantees are useful, they only describe the regret distribution at a single point and require the user to choose $\delta$ in advance. By contrast, in many safety-critical applications, such as healthcare decision making, robotic control in human environments, and financial risk management Inamdar et al. (2024), one would like a more detailed characterization of the entire tail distribution of regret, that is, explicit bounds on $\mathbb{P}(R_{K}\geq x)$ that hold for all thresholds $x\geq\mathbb{E}[R_{K}]$.
In this paper, we take a step towards closing this gap by studying the concentration of regret for a UCBVI-type algorithm in finite-horizon MDPs. We show that the tail of the cumulative regret, $\mathbb{P}(R_{K}\geq x)$, admits an explicit, non-asymptotic sub-Gaussian or sub-Weibull bound, depending on the value of $x$ and the instance-dependent $Q^{\star}$-gap parameter.
Our contributions are as follows:
  • We study the tail behavior of the cumulative regret $R_{K}$ for a UCBVI-type algorithm in finite-horizon tabular MDPs under a global $Q^{\star}$-gap, and derive explicit, instance-dependent sub-Gaussian and sub-Weibull tail bounds for $\mathbb{P}(R_{K}\geq x)$. Our analysis treats two families of exploration bonuses: a $K$-dependent schedule whose bonuses depend on the total number of episodes $K$, and a $K$-independent schedule whose bonuses depend only on the current episode index.
  • For the $K$-dependent scheme, we show that the regret exhibits a sub-Gaussian tail on an intermediate regime of $x$ and a sub-Weibull tail beyond $x>\tilde{\mathcal{O}}\!\big(H^{5/2}K^{(1+\alpha)/2}\big)$, where $\alpha\in[0,1]$ is a tuning parameter.
  • For the $K$-independent scheme, we obtain a similar two-regime behavior, where the transition to sub-Weibull happens at $\tilde{\mathcal{O}}\!\big(H^{(5-\alpha)/(2-\alpha)}K^{1/(2-\alpha)}\big)$.
  • We further show that the expected regret is $\mathbb{E}[R_{K}]\;\leq\;\tilde{\mathcal{O}}\!\Big(H^{6}SA(S+K^{\alpha})/\mathrm{gap}^{*}\Big)$, where $\mathrm{gap}^{*}>0$ is the global $Q^{\star}$-gap. The tuning parameter $\alpha$ balances the expected regret and the range over which the regret exhibits a sub-Gaussian tail.
Our results provide the first characterization of the tail of the regret for a standard optimistic algorithm in tabular episodic RL, and they suggest that the entire distribution of the regret—not just its expectation or a single high-probability quantile—can be controlled in an instance-dependent manner determined by the $Q^{\star}$-gap.
2 Related Work
Worst-case regret.
The modern theory of online RL in finite MDPs starts from worst-case regret guarantees under the optimism in the face of uncertainty principle. In the average-reward setting, the UCRL2 algorithm of Jaksch et al. (2010) achieves near-minimax regret of order $\widetilde{\mathcal{O}}\!\big(DS\sqrt{AK}\big)$, where $D$ is the diameter of the MDP, and lower bound $\widetilde{\mathcal{O}}\!\big(\sqrt{DSAK}\big)$ shows that the dependence on $(D,S,A,K)$ is almost optimal. Subsequent refinements sharpened the constants and logarithmic factors, for instance by tightening the confidence sets around the transition kernel via empirical Bernstein-type bonuses (Bourel et al., 2020; Fruit et al., 2020).
In the episodic finite-horizon setting considered in this paper, the optimal worst-case scaling was established by the UCBVI algorithm of Azar et al. (2017), which attains regret $\widetilde{\mathcal{O}}\!\big(H\sqrt{SAK}+H\sqrt{KH}\big)$ and derive information-theoretic lower bounds of $\widetilde{\mathcal{O}}\!\big(H\sqrt{SAK}\big)$. A number of works have since refined or extended this result: Efroni et al. (2019) proposed the EULER algorithm, which yields worst-case bounds with improved dependence on the horizon; see also the recent minimax analyses and extensions to factored MDPs, e.g., Tian et al. (2020). These results, however, are primarily worst-case and focus on bounding either the expectation or a single high-probability quantile of the regret, without characterizing the entire tail behavior.
Instance-dependent and gap-dependent analysis.
Motivated by the classical gap-dependent theory for stochastic bandits (see, e.g., the survey of Bubeck and Cesa-Bianchi (2012)), there has been a surge of interest in instance-dependent regret bounds for tabular MDPs. Simchowitz and Jamieson (2019) derived the first non-asymptotic gap-dependent regret bounds for model-based tabular RL, showing that the regret can scale polylogarithmically in $K$ (up to polynomial factors in $(S,A,H)$ and $1/\mathrm{gap}^{*}$). Zanette and Brunskill (2019) introduces the environmental norm, a variance-like quantity that controls the difficulty of exploration, and designed an optimistic algorithm whose regret depends on this norm rather than only on worst-case parameters. More recently, Dann et al. (2021) developed improved instance-dependent bounds that replace the global value-function gaps with more local notions tailored to the subset of states visited by optimal policies, together with new information-theoretic lower bounds showing the necessity of these refined gap notions. On the model-free side, Yang et al. (2021) showed that optimistic $Q$-learning enjoys logarithmic regret in episodic tabular MDPs under a global suboptimality-gap on $Q^{\star}$, thus establishing that simple temporal-difference methods can match the best-possible gap-dependent rates of model-based algorithms. Xu et al. (2021) proposed an adaptive multi-step bootstrapping scheme that yields tighter instance-dependent bounds.
Under the linear MDP assumption with function approximation, He et al. (2021) establish instance-dependent logarithmic regret for LSVI-UCB and UCRL-VTR in linear and linear-mixture MDPs. Velegkas et al. (2022) further refine this picture by designing algorithms that incur only a bounded number of policy switches. Most recently, Zhao et al. (2025) extend these guarantees to adversarial environments.
Tail behavior and risk-sensitive objectives.
A recent line of work, starting from Fan and Glynn (2024), shows that bandit algorithms that are information-theoretically optimized for Lai–Robbins optimal expected regret can exhibit extremely heavy-tailed regret distributions: the regret tail behaves like a truncated Cauchy, and higher moments grow polynomially in the horizon. Building on this observation, Simchi-Levi et al. (2022, 2023) design policies that explicitly trade off optimality and tail risk, achieving worst-case optimal (or near-optimal) expected regret while simultaneously enforcing exponentially light tails for large-regret events and characterizing sharp Pareto frontiers between expected regret and tail probabilities.
In contrast, RL work that explicitly targets tail or risk-sensitive criteria usually changes the performance objective rather than analyzing the tail of classical regret. For example, there is a growing literature on risk-sensitive or distributional RL that optimizes mean-variance, CVaR, or more general optimized certainty equivalent (OCE) criteria in MDPs; see, e.g., Tamar et al. (2015); Chow et al. (2015); Xu et al. (2023) and references therein. These works derive sample-complexity or regret guarantees for the risk-sensitive value of a policy, not for the tail of the standard regret with respect to the risk-neutral optimal policy.
Our work is complementary to this risk-sensitive line: we keep the classical regret objective but ask for an instance-dependent characterization of its tail distribution. In particular, given a global $Q^{\star}$-gap, we show that a UCBVI-type algorithm enjoys sub-Gaussian and sub-Weibull tail bounds for the cumulative regret, together with instance-dependent control of its expectation. To the best of our knowledge, this is the first result that treats regret concentration in episodic RL.
3 Problem Formulation
We consider an MDP $(\mathcal{S},\mathcal{A},\bm{P}^{*},r,H)$ with finite horizon $H$, where $\mathcal{S}$ and $\mathcal{A}$ are finite state and action spaces with $S\coloneqq|\mathcal{S}|$ and $A\coloneqq|\mathcal{A}|$, the transition kernel is $\bm{P}^{*}=\{P_{h}^{*}\}_{h=0}^{H-2}$, and the reward function is $r=\{r_{h}\}_{h=0}^{H-1}$, where each stage-$h$ reward is given by $r_{h}:\mathcal{S}\times\mathcal{A}\to[0,1]$.
A deterministic Markov policy is a collection $\bm{\pi}=\{\pi_{h}\}_{h=0}^{H-1}$ with $\pi_{h}:\mathcal{S}\to\mathcal{A}$. Each episode starts at a fixed initial state $s_{0}\in\mathcal{S}$ and lasts exactly $H$ steps. Given any state $s_{h}$ at horizon $h$, following policy $\bm{\pi}$, we take the action $a_{h}=\pi_{h}(s_{h})$ and transition to the next random state $s_{h+1}\sim P_{h}^{*}(\cdot\mid s_{h},a_{h})$. The value and action-value functions are defined as
$V_{h}^{\bm{\pi}}(s)$ $\;=\;\mathbb{E}_{\bm{\pi},\bm{P}^{*}}\!\Big[\sum_{t=h}^{H-1}r_{t}(s_{t},a_{t})\,\Big|\,s_{h}=s\Big],$
$Q_{h}^{\bm{\pi}}(s,a)$ $\;=\;r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}}P_{h}^{*}(s^{\prime}\mid s,a)\,V_{h+1}^{\bm{\pi}}(s^{\prime}),$
with the convention $V_{H}^{\bm{\pi}}\equiv 0$. Here $\mathbb{E}_{\bm{\pi},\bm{P}^{*}}[\cdot]$ denotes expectation with respect to policy $\bm{\pi}$ and transition kernel $\bm{P}^{*}$. The optimal value and action-value functions are
$V_{h}^{*}(s)\coloneqq\max_{\bm{\pi}}V_{h}^{\bm{\pi}}(s),\qquad Q_{h}^{*}(s,a)\coloneqq r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}}P_{h}^{*}(s^{\prime}\mid s,a)\,V_{h+1}^{*}(s^{\prime}),$
and the set of optimal actions at $(s,h)$ is
$\mathcal{A}_{h}^{*}(s)\coloneqq\arg\max_{a\in\mathcal{A}}Q_{h}^{*}(s,a).$
We write $\mathbf{\Pi}^{*}$ for the set of optimal deterministic Markov policies. Our analysis is instance-dependent and exploits a global $Q^{*}$-gap:
$\mathrm{gap}_{h}(s,a)\;$ $\coloneqq\;\Big(Q_{h}^{*}(s,\pi_{h}^{*}(s))-Q_{h}^{*}(s,a)\Big),\qquad\forall(s,h)\in\mathcal{S}\times[H],$
$\mathrm{gap}^{*}\;$ $\coloneqq\;\min_{\begin{subarray}{c}(s,h)\in\mathcal{S}\times[H]\\ a\notin\mathcal{A}_{h}^{*}(s)\end{subarray}}\mathrm{gap}_{h}(s,a),$ (1)
where $\bm{\pi}^{*}\in\mathbf{\Pi}^{*}$ is any optimal Markov policy.
Interaction proceeds over $K$ episodes indexed by $k\in[K]$ [^1] Throughout, $n]=\{0,1,\dots,n-1\}$ and $n_{1},n_{2}]=\{n_{1},n_{1}+1,\dots,n_{2}-1\}$.. At the start of episode $k$, the learner commits to a Markov policy $\bm{\pi}^{k}=\{\pi_{h}^{k}\}_{h=0}^{H-1}$, then observes a trajectory
$s_{0},\ a_{0}^{k},\ s_{1}^{k},\ a_{1}^{k},\ \dots,\ s_{H-1}^{k},\ a_{H-1}^{k},$
where $s_{0}^{k}=s_{0}$, $a_{h}^{k}=\pi_{h}^{k}(s_{h}^{k})$, and $s_{h+1}^{k}\sim P_{h}^{*}(\cdot\mid s_{h}^{k},a_{h}^{k})$ for all $h\in\{0,\dots,H-2\}$. The performance of the learner is measured by the cumulative regret
$R_{K}\;=\;\sum_{k=0}^{K-1}\Big(V_{0}^{\bm{\pi}^{*}}(s_{0})-V_{0}^{\bm{\pi}^{k}}(s_{0})\Big),$
where $\bm{\pi}^{*}\in\mathbf{\Pi}^{*}$ is any optimal policy. Our goal is to develop an algorithm that attains low expected cumulative regret and, simultaneously, exhibits a desirable tail distribution for $R_{K}$.
4 Main Results
In this section, we present our main results. We study the tail behavior of the regret for a UCBVI-type algorithm in the model-based setting. We analyze a UCBVI-type algorithm that uses the following bonus function:
$b^{k}_{h}(n)=\begin{cases}(H-h-1)\sqrt{\dfrac{0.5S\ln 2+\mu K^{\alpha}}{n}},&\text{{KD}},\\[8.61108pt] (H-h-1)\sqrt{\dfrac{0.5S\ln 2+\mu(k+1)^{\alpha}}{n}},&\text{{KI}},\end{cases}\qquad b^{k}_{h}(0)\coloneqq\infty.$
Here $k\in\{0,\dots,K-1\}$ denotes the episode index, $n\geq 1$ denotes the number of visits to the corresponding state–action pair, and $\alpha\in[0,1]$ and $\mu>0$ are tunable parameters. KD refers to the bonus used in the $K$-dependent algorithm, and KI refers to the bonus used in the $K$-independent algorithm.
The term $K^{\alpha}$ in the definition of $b_{h}^{k}(\cdot)$ plays the role of a global exploration budget that directly controls the aggressiveness of the bonuses. In the original minimax UCBVI analysis of Azar et al. (2017), as well as in the subsequent instance-dependent analysis of Simchowitz and Jamieson (2019), the exploration terms scale only logarithmically with the time horizon, roughly as $\sqrt{{\log(KSAH)}/{n}}$, which is tailored to control of the expected regret. In contrast, our bonuses scale as $\sqrt{(S\ln 2+\mu K^{\alpha})/n}$ (or $\sqrt{(S\ln 2+\mu(k+1)^{\alpha})/n}$ in the $K$-independent case), so that the uncertainty radius around the empirical model grows polynomially in $K$. This more conservative choice incentivizes sustained exploration even in later episodes: when the algorithm has visited a state–action pair only moderately many times, the bonus remains substantial and pushes the optimistic value estimates upward. From a tail perspective, this is precisely what allows us to derive nontrivial bounds on $\mathbb{P}(R_{K}\geq x)$ for large thresholds $x$, at the price of a controlled polynomial dependence on $K$ inside the bonus which will affect the expected regret.
Algorithm 1 Optimistic Value Iteration
1:Input: $H,\mathcal{S},\mathcal{A}$, rewards $r_{h}:\mathcal{S}\times\mathcal{A}\to[0,1]$, initial state $s_{0}$, bonus schedules $\{b^{k}_{h}(\cdot)\}$.
2:Init: For all $(s,a,h)$, $n_{h}^{0}(s,a)=0$ and $\hat{P}_{h}^{0}(\cdot\mid s,a)=\tfrac{1}{S}$, and $\hat{P}_{H-1}^{k}\equiv 0$ and all $k$.
3:for $k=0,1,\dots,K-1$ do
4:  Planning: Set $V_{H}^{k}\equiv 0$. For $h=H-1,\dots,0$ and each $s$:
$Q_{h}^{k}(s,a)$ $\coloneqq\min\left\{r_{h}(s,a)+\sum_{s^{\prime}}\hat{P}_{h}^{k}(s^{\prime}\mid s,a)\,V_{h+1}^{k}(s^{\prime})+b_{h}^{k}(n_{h}^{k}(s,a)),H-h\right\},\quad\forall a\in\mathcal{A},$
$V_{h}^{k}(s)$ $\coloneqq\max_{a\in\mathcal{A}}Q_{h}^{k}(s,a),\qquad\pi_{h}^{k}(s)\in\arg\max_{a}Q_{h}^{k}(s,a).$
5:  Execution: Roll out $\bm{\pi}^{k}=\{\pi_{h}^{k}\}_{h=0}^{H-1}$ from $s_{0}$ under $\bm{P}^{*}$:
$a_{h}^{k}=\pi_{h}^{k}(s_{h}^{k})\quad\text{and}\quad s_{h+1}^{k}\sim P_{h}^{*}(\cdot\mid s_{h}^{k},a_{h}^{k}),\qquad\forall h\in[H-1],\qquad a_{H-1}^{k}=\pi_{H-1}^{k}(s_{H-1}^{k}).$
6:  Update: For all $(s,a,h)$:
$n_{h}^{k+1}(s,a,s^{\prime})=\sum_{i=0}^{k}\mathbbm{1}\{(s_{h}^{i},a_{h}^{i},s_{h+1}^{i})=(s,a,s^{\prime})\},\quad n_{h}^{k+1}(s,a)=\sum_{s^{\prime}}n_{h}^{k+1}(s,a,s^{\prime}),$
$\hat{P}_{h}^{k+1}(s^{\prime}\mid s,a)=\begin{cases}\dfrac{n_{h}^{k+1}(s,a,s^{\prime})}{n_{h}^{k+1}(s,a)},&n_{h}^{k+1}(s,a)\geq 1,\\[4.30554pt] \dfrac{1}{S},&n_{h}^{k+1}(s,a)=0.\end{cases}$
7:end for
The UCBVI algorithm with the above bonus function is presented in Algorithm 1. The input consists of the finite-horizon MDP parameters $(H,\mathcal{S},\mathcal{A},r)$, the initial state $s_{0}$, and the family of bonus schedules $\{b_{h}^{k}(\cdot)\}$. In the initialization step, the algorithm sets all visit counts $n_{h}^{0}(s,a)$ to zero and uses the uniform distribution $\hat{P}_{h}^{0}(\cdot\mid s,a)=\tfrac{1}{S}$ as an initial model of the dynamics.
At the beginning of each episode $k$, the algorithm performs a planning phase: it runs a backward dynamic-programming pass from h=H1h=H-1 down to $h=0$, maintaining an optimistic value function $V_{h}^{k}$ and $Q$-function $Q_{h}^{k}$. For each state-action pair $(s,a)$ at horizon $h$, it forms the optimistic estimate
$Q_{h}^{k}(s,a)=\min\left\{r_{h}(s,a)+\sum_{s^{\prime}}\hat{P}_{h}^{k}(s^{\prime}\mid s,a)\,V_{h+1}^{k}(s^{\prime})+b_{h}^{k}\big(n_{h}^{k}(s,a)\big),\,H-h\right\},$
where the bonus $b_{h}^{k}(n_{h}^{k}(s,a))$ inflates the value of actions whose transition probabilities are still poorly estimated. The value function $V_{h}^{k}(s)$ is then defined using the greedy policy $\pi_{h}^{k}(s)$ which selects any maximizer of $Q_{h}^{k}(s,\cdot)$. This can be viewed as solving the backward dynamic equations with the estimated transition kernel $\hat{\bm{P}}^{k}$ and the modified reward function $r_{h}(\cdot,\cdot)+b_{h}^{k}\big(n_{h}^{k}(\cdot,\cdot)\big).$
In the execution phase, the agent rolls out the greedy policy $\pi^{k}$ for one episode starting from $s_{0}$, interacting with the true MDP $\bm{P}^{*}$ and collecting a trajectory $\{(s_{h}^{k},a_{h}^{k},s_{h+1}^{k})\}_{h=0}^{H-1}$. Finally, in the update phase, the algorithm increments the visit counts $n_{h}^{k+1}(s,a,s^{\prime})$ and $n_{h}^{k+1}(s,a)$ along the realized trajectory and updates the empirical transition kernel $\hat{P}_{h}^{k+1}(\cdot\mid s,a)$ by normalized counts. The procedure then repeats with the updated model and bonuses in the next episode.
Theorem 1 characterizes the tail behavior of the regret of Algorithm 1. In this theorem, we use the notation $(x)_{+}\coloneqq\max\{x,0\}$.
Theorem 1.
Fix $\alpha\in[0,1]$ and $\mu>0$. For all $\gamma\in[0,1]$, the tail of the cumulative regret $R_{K}$ of Algorithm 1 satisfies
$\mathbb{P}(R_{K}\geq x)\ \leq\ \exp\!\left(-\frac{\big(x-H\lceil K^{\gamma}\rceil-\,m_{K}^{\mathrm{mb}}\big)_{+}^{\,2}}{50H^{5}K}\right)\;+\;\delta_{K}^{\mathrm{mb}}(\gamma),\qquad\forall x\geq H\lceil K^{\gamma}\rceil,$
where
$m_{K}^{\mathrm{mb}}\;\coloneqq\frac{32H^{6}SA(0.5S\ln 2+\mu K^{\alpha})}{\mathrm{gap}^{*}},$ (2)
and
$\delta_{K}^{\mathrm{mb}}(\gamma)=SAHK\begin{cases}\displaystyle\exp\left(-2\mu K^{\alpha}\right),&\text{{KD}},\\[8.61108pt] \displaystyle\exp\left(-2\mu K^{\gamma\alpha}\right),&\text{{KI}}.\end{cases}$ (3)
Theorem 1 shows that, under a global $Q^{\star}$-gap, as defined in (1), the tail of the cumulative regret $R_{K}$ admits a decomposition into a sub-Gaussian component and a residual term $\delta_{K}^{\mathrm{mb}}(\gamma)$. The quantity $m_{K}^{\mathrm{mb}}$ is the baseline instance-dependent level of regret around which the distribution concentrates: it aggregates the bonuses incurred before all suboptimal actions have been sufficiently explored and eliminated, and it is of order
$m_{K}^{\mathrm{mb}}\;\approx\;\mathcal{O}\!\big(H^{6}SA(S+K^{\alpha})/\mathrm{gap}^{*}\big)$
up to logarithmic factors. For thresholds $x$ that are moderately larger than $m_{K}^{\mathrm{mb}}$, the bound $\mathbb{P}(R_{K}\geq x)$ is of sub-Gaussian form with variance proxy $25H^{5}K$, reflecting the Martingale structure of the regret fluctuations after the main exploration phase.
Proposition 1.
Under the same setting as in Theorem 1, we have
$\mathbb{E}[R_{K}]\ \leq\ H\lceil K^{\gamma}\rceil+m_{K}^{\mathrm{mb}}+6KH^{3}\delta_{K}^{\mathrm{mb}}(\gamma),$
where $m_{K}^{\mathrm{mb}}$ is given in (2) and $\delta_{K}^{\mathrm{mb}}(\gamma)$ in (3).
Proposition 1 obtains an instance-dependent upper bound on the expected regret. The first two terms, $m_{K}^{\mathrm{mb}}+H\lceil K^{\gamma}\rceil$, capture the contribution of the baseline instance-dependent level of regret and the initial “burn-in” phase. The final term, proportional to $6KH^{3}\delta_{K}^{\mathrm{mb}}(\gamma)$, accounts for the rare events in which the empirical model falls outside the “good” set at some episodes. In such events the estimated transition kernel $\hat{\mathbb{P}}^{k}$ departs substantially from the true kernel $\mathbb{P}^{*}$ for some $k\in[K]$. Consequently, we cannot ensure that $V^{k}_{0}(s_{0})\geq V^{*}_{0}(s_{0})$ uniformly over $k$, i.e., the optimism property may fail. Thanks to the exponential decay of $\delta_{K}^{\mathrm{mb}}(\gamma)$ in $K^{\alpha}$ (or $K^{\gamma\alpha}$), this residual contribution is negligible: for any fixed $\alpha>0$ and suitable choice of the constant $\mu>0$, the product $K\delta_{K}^{\mathrm{mb}}(\gamma)$ vanishes.
The tuning parameter $\alpha\in[0,1]$ controls the tradeoff between the expected regret and the decay rate of the residual term $\delta_{K}^{\mathrm{mb}}(\gamma)$. Larger values of $\alpha$ make the bonuses more aggressive, which increases the upper bound on the $\mathbb{E}[R_{K}]$ but also causes $\delta_{K}^{\mathrm{mb}}(\gamma)$ to decay faster with $K$ (either as $\exp(-2\mu K^{\alpha})$ or as $\exp(-2\mu K^{\gamma\alpha})$ depending on the bonus schedule). In this sense, $\alpha$ interpolates between prioritizing a smaller expected regret ($\alpha$ small) and enforcing lighter tails for very large deviations ($\alpha$ closer to one). By contrast, the parameter $\gamma\in[0,1]$ is not a tuning knob of the algorithm: it appears only in the analysis and controls the length of the initial burn-in phase $K^{\gamma}$ during which we make no attempt to exploit tail concentration. In Corollary 1, we choose $\gamma$ as a function of $x$ to simplify the resulting tail bound into a clean two-regime description.
Corollary 1.
Fix $\alpha\in[0,1]$. For the KD algorithm, we have
$\mathbb{P}(R_{K}\geq x)\leq\begin{cases}\exp\left(-\tilde{\mathcal{O}}\left(\frac{x^{2}}{H^{5}K}\right)\right)&\text{if }\mathcal{O}(H^{6}SA(S+K^{\alpha})/\mathrm{gap}^{*})\leq x\leq\tilde{\mathcal{O}}(H^{5/2}K^{(1+\alpha)/2})\\ \exp\left(-\tilde{\mathcal{O}}\left(K^{\alpha}\right)\right)&\text{if }x\geq\tilde{\mathcal{O}}(H^{5/2}K^{(1+\alpha)/2})\end{cases}$
For KI algorithm, we have
$\mathbb{P}(R_{K}\geq x)\leq\begin{cases}\exp\left(-\tilde{\mathcal{O}}\left(\frac{x^{2}}{H^{5}K}\right)\right)&\text{if }\mathcal{O}(H^{6}SA(S+K^{\alpha})/\mathrm{gap}^{*})\leq x\leq\tilde{\mathcal{O}}(H^{(5-\alpha)/(2-\alpha)}K^{1/(2-\alpha)})\\ \exp\left(-\tilde{\mathcal{O}}\left(\frac{x^{\alpha}}{H^{\alpha}}\right)\right)&\text{if }x\geq\tilde{\mathcal{O}}(H^{(5-\alpha)/(2-\alpha)}K^{1/(2-\alpha)})\end{cases}$
Furthermore, $\mathbb{E}[R_{K}]\leq\mathcal{O}(H^{6}SA(S+K^{\alpha})/\mathrm{gap}^{*})$.
Proof of Corollary 1.
The proof follows by setting $\gamma=0$ for KD algorithm, and setting $\gamma=\log_{K}(x/2H)$ for KI algorithm. ∎
It is natural to compare Corollary 1 with the existing gap-dependent bounds for tabular MDPs. For model-based algorithms, Simchowitz and Jamieson (2019) show that, given a global $Q^{\star}$-gap, for any fixed confidence level $\delta\in(0,1)$, the cumulative regret satisfies the following high-probability bound:
$\mathbb{P}\left(R_{K}\geq\tilde{\mathcal{O}}\!\Big(H^{4}SA(S\vee H)+\frac{H^{3}SA}{\mathrm{gap}^{*}}\log(K/\delta)\Big)\right)\;\leq\;\delta.$
However, they do not characterize the tail behavior $\mathbb{P}(R_{K}\geq x)$ for all $x$. Setting $\delta=1/K$, this high probability bound implies the expected regret that scale polylogarithmically in $K$, with bounds of the form
$\mathbb{E}[R_{K}]\;\leq\;\tilde{\mathcal{O}}\!\Big(\frac{H^{3}SA}{\mathrm{gap}^{*}}\log K\Big).$
Similar instance-dependent rates have been obtained for optimistic $Q$-learning and related model-free methods (Yang et al., 2021; Velegkas et al., 2022). All the algorithms presented in these works depend on the specific choice of confidence level $\delta\in(0,1)$.
In contrast, our results are instance-dependent in a comparable regime, but they are organized around the entire tail rather than a single quantile. Instead of working at a fixed confidence level $\delta$ and producing a bound, Corollary 1 describes how the entire curve $x\mapsto\mathbb{P}(R_{K}\geq x)$ behaves: for $x$ above the instance-dependent baseline $m_{K}^{\mathrm{mb}}$ we obtain a sub-Gaussian tail
$\mathbb{P}(R_{K}\geq x)\;\leq\;\exp\!\Big(-\tilde{\mathcal{O}}\big(x^{2}/(H^{5}K)\big)\Big),$
up to a transition scale that depends on $(H,K,\alpha)$, and beyond that scale the tail becomes sub-Weibull, with decay of order $\exp\{-\tilde{\mathcal{O}}(K^{\alpha})\}$ in the KD case or $\exp\{-\tilde{\mathcal{O}}(x^{\alpha}/H^{\alpha})\}$ in the KI case.
5 Proof Analysis
5.1 Proof of Theorem 1
Proof of Theorem 1.
Recall that $\bm{\pi}^{*}=\{\pi^{*}_{h}\}_{h=0}^{H-1}\in{\mathbf{\Pi}}^{*}$, where $\mathbf{\Pi}$$^{*}$ denotes the set of optimal policies for the transition kernel $\bm{P}^{*}=\{P^{*}_{h}\}_{h=0}^{H-2}$. At episode $k$, the algorithm computes a greedy policy $\pi^{k}=\{\pi^{k}_{h}\}_{h=0}^{H-1}$ using an optimistic estimate of the optimum value function $\{V^{*}_{h}\}_{h=0}^{H-1}$ based on data collected up to episode $k$. Recall that $R_{K}$ denote the cumulative regret incurred up to episode K1K-1. Then, writing $s_{0}^{k}\equiv s_{0}$, we have:
$R_{K}$ $=\sum_{k=0}^{K-1}\left(V_{0}^{*}(s_{0}^{k})-V_{0}^{\bm{\pi}^{k}}(s_{0}^{k})\right)=\underbrace{\sum_{k=0}^{\lceil K^{\gamma}\rceil-1}\left(V_{0}^{*}(s_{0}^{k})-V_{0}^{\bm{\pi}^{k}}(s_{0}^{k})\right)}_{R_{K}^{(0)}}+\underbrace{\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left(V_{0}^{*}(s_{0}^{k})-V_{0}^{\bm{\pi}^{k}}(s_{0}^{k})\right)}_{R_{K}^{(1)}}.$
Here, $R_{K}^{(0)}\leq\lceil K^{\gamma}\rceil H$ denotes the initial regret incurred during episodes with insufficient data. We continue by analyzing $R_{K}^{(1)}$. Throughout this proof, we define the following clipping function:
$\mathrm{clip}(x)\;=\;\begin{cases}0,&\text{if }x<\mathrm{gap}^{*},\\[4.0pt] x,&\text{otherwise.}\end{cases}$
Note that for any non negative numbers $a_{1},a_{2},\dots,a_{H}$, the $\mathrm{clip}$ function satisfies
$\mathrm{clip}\left(\sum_{i=1}^{H}a_{i}\right)\leq\sum_{i=1}^{H}\mathrm{clip}(Ha_{i}).$ (4)
The regret associated with any episode $k\geq\left\lceil K^{\gamma}\right\rceil$ is given by
$V_{0}^{*}(s_{0}^{k})-V_{0}^{\bm{\pi}^{k}}(s_{0}^{k})$ $=V_{0}^{*}(s_{0}^{k})-Q_{0}^{*}(s_{0}^{k},a_{0}^{k})+(Q_{0}^{*}(s_{0}^{k},a_{0}^{k})-Q_{0}^{\bm{\pi}^{k}}(s_{0}^{k},a_{0}^{k}))$
$=\mathrm{gap}_{0}(s_{0}^{k},a_{0}^{k})+\mathbb{E}\left[V_{1}^{*}(s_{1}^{k})-V_{1}^{\bm{\pi}^{k}}(s_{1}^{k})\big|\mathcal{F}_{0}^{k}\right]$
where $\mathcal{F}_{h}^{k}=\sigma(\{s_{0}^{0},a_{0}^{0},s_{1}^{0},a_{1}^{0},\ldots,s_{H-1}^{0},a_{H-1}^{0},s_{0}^{1},a_{0}^{1},\ldots,s_{0}^{k},a_{0}^{k},\ldots,s_{h}^{k},a_{h}^{k}\})$ represents the history of states and actions observed by the algorithm up to episode $k$ and step $h$, and $a_{h}^{k}=\pi_{h}^{k}(s_{h}^{k})$. Define the value difference $\Delta_{h}^{k}(s)\coloneqq V_{h}^{*}(s)-V_{h}^{\bm{\pi}^{k}}(s)$. Following a similar argument as above, for all $h\in[H]$, we have
$\Delta^{k}_{h}(s_{h}^{k})$ $=\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})+\mathbb{E}\left[\Delta^{k}_{h+1}(s_{h+1}^{k})\big|\mathcal{F}_{h}^{k}\right]$
$=\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})+\Delta^{k}_{h+1}(s_{h+1}^{k})-\eta^{k}_{h+1}(s_{h+1}^{k}),$
where the sequence $\{\eta^{k}_{h}(s_{h}^{k})\}_{h=1,k=0}^{H-1,K-1}$ defined as
$\eta^{k}_{h}(s_{h}^{k})=\Delta^{k}_{h}(s_{h}^{k})-\mathbb{E}\left[\Delta^{k}_{h}(s_{h}^{k})\big|\mathcal{F}_{h-1}^{k}\right]$
is a Martingale difference sequence with the filtration $\{\mathcal{F}_{h-1}^{k}\}_{h=1,k=0}^{H-1,K-1}$. For the sake of notational simplicity, we define $\eta^{k}_{H}\equiv 0$ for all $k\in[K]$. Using a telescopic sum, noting that $\Delta^{k}_{H}\equiv 0$, we obtain
$\Delta^{k}_{0}(s_{0}^{k})$ $=\sum_{h=0}^{H-1}\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})$
$=\sum_{h=0}^{H-1}\mathrm{clip}\left(\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})\right)-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})$ (5)
where the last equality follows by the fact that $\mathrm{gap}_{h}(s,a)\geq\mathrm{gap}^{*}$ if $a\notin\mathcal{A}^{*}_{h}(s)$ and $\mathrm{gap}_{h}(s,a)=0$ otherwise.
Note that $\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})=V_{h}^{*}(s_{h}^{k})-Q_{h}^{*}(s_{h}^{k},a_{h}^{k})\leq V_{h}^{*}(s_{h}^{k})-V_{h}^{k}(s_{h}^{k})+V_{h}^{k}(s_{h}^{k})-V_{h}^{\bm{\pi}^{k}}(s_{h}^{k})$. The bonus function $b_{h}^{k}(\cdot)$ is chosen so that $V_{h}^{k}$ provides an optimistic estimate of $V_{h}^{*}$, as long as the estimated transition kernel $\hat{P}_{h^{\prime}}^{k}$ and the true transition kernel $P_{h^{\prime}}^{*}$ remain close for all horizons $h^{\prime}\geq h$. This combined with the monotonicity of the $\mathrm{clip}(\cdot)$ function provides an upper bound for the regret at episode $k\geq\lceil K^{\gamma}\rceil$. We define the “good event at episode kk” as
$\mathcal{G}^{k}=$ $\bigcap_{h=0}^{H-2}\bigcap_{s,a}\left\{\|P_{h}^{*}(\cdot|s,a)-\hat{P}_{h}^{k}(\cdot|s,a)\|_{1}\leq\frac{2b_{h}^{k}(n_{h}^{k}(s,a))}{H-h-1}\right\}.$
Under the event $\mathcal{G}^{k}$, we claim that $Q_{h}^{k}(s,a)\geq Q_{h}^{*}(s,a)$ for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$. We establish this by induction on the horizon. At the terminal horizon, we have $Q_{H-1}^{k}(s,a)=Q_{H-1}^{*}(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. Assuming $Q_{h+1}^{k}(s,a)\geq Q_{h+1}^{*}(s,a)$ for all $(s,a)$, we now show that the same inequality holds for horizon $h$.
$Q_{h}^{k}(s,a)-Q_{h}^{*}(s,a)=\min\Bigg\{$ $\sum_{s^{\prime}}\hat{P}_{h}^{k}(s^{\prime}|s,a)\big(V_{h+1}^{k}(s^{\prime})-V_{h+1}^{*}(s^{\prime})\big)$
$\quad+\sum_{s^{\prime}}\big(\hat{P}_{h}^{k}(s^{\prime}|s,a)-P_{h}^{*}(s^{\prime}|s,a)\big)V_{h+1}^{*}(s^{\prime})+b_{h}^{k}\big(n_{h}^{k}(s,a)\big),$
$\quad H-h-Q_{h}^{*}(s,a)\Bigg\}.$
The first term is nonnegative by induction hypothesis, while the second term is bounded as follows:
$\sum_{s^{\prime}}\!\left(\hat{P}_{h}^{k}(s^{\prime}|s,a)-P_{h}^{*}(s^{\prime}|s,a)\right)V_{h+1}^{*}(s^{\prime})$
$\qquad=\sum_{s^{\prime}}\!\left(\hat{P}_{h}^{k}(s^{\prime}|s,a)-P_{h}^{*}(s^{\prime}|s,a)\right)\left(V_{h+1}^{*}(s^{\prime})-c\mathbf{1}\right)$
$\qquad\geq-\|\hat{P}_{h}^{k}(\cdot|s,a)-P_{h}^{*}(\cdot|s,a)\|_{1}\,\|V_{h+1}^{*}-c\mathbf{1}\|_{\infty}$ (Hölder’s inequality)
$\qquad=-\tfrac{1}{2}\,\mathrm{span}\big(V_{h+1}^{*}\big)\,\|\hat{P}_{h}^{k}(\cdot|s,a)-P_{h}^{*}(\cdot|s,a)\|_{1}$
$\qquad\geq-b_{h}^{k}(n_{h}^{k}(s,a)),$ (by definition of ${\mathcal{G}}^{k}$)
where $c=\tfrac{1}{2}\big(\max_{x}V_{h+1}^{*}(x)+\min_{x}V_{h+1}^{*}(x)\big),$ $\mathrm{span}(f)=\max_{x}f(x)-\min_{x}f(x)$, and $\mathbf{1}$ denotes the all-ones vector. Hence, under $\bar{\mathcal{G}}^{K}$, taking maxima and recursing yields
$Q_{h}^{k}(s,a)\geq Q_{h}^{*}(s,a),\qquad V_{h}^{k}(s)\geq V_{h}^{*}(s),$ (6)
for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$.
Defining $\bar{\mathcal{G}}^{K}\coloneqq\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}^{k}$, we next bound the probability of the bad event $\left(\bar{\mathcal{G}}^{K}\right)^{\mathrm{c}}$:
$\mathbb{P}\left(\left(\bar{\mathcal{G}}^{K}\right)^{\mathrm{c}}\right)$
$=\mathbb{P}\left(\exists s,a,h,k\in[\lceil K^{\gamma}\rceil,K]:\|P_{h}^{*}(\cdot|s,a)-\hat{P}_{h}^{k}(\cdot|s,a)\|_{1}>\frac{2b_{h}^{k}(n_{h}^{k}(s,a))}{H-h-1}\right)$
$\leq\sum_{s,a,h}\mathbb{P}\left(\exists k\in[\lceil K^{\gamma}\rceil,K]:\sum_{s^{\prime}}\left|P_{h}^{*}(s^{\prime}|s,a)-\frac{n_{h}^{k}(s,a,s^{\prime})}{n_{h}^{k}(s,a)}\right|>\frac{2b_{h}^{k}(n_{h}^{k}(s,a))}{H-h-1}\right)$ (union bound)
$=\sum_{s,a,h}\mathbb{P}\left(\exists k\in[\lceil K^{\gamma}\rceil,K]:\sum_{s^{\prime}}\left|P_{h}^{*}(s^{\prime}|s,a)-\frac{\sum_{i=0}^{k-1}\mathbbm{1}\{(s^{i}_{h},a^{i}_{h},s^{i}_{h+1})=(s,a,s^{\prime})\}}{n_{h}^{k}(s,a)}\right|>\frac{2b_{h}^{k}(n_{h}^{k}(s,a))}{H-h-1}\right)$
$\overset{(a)}{\leq}\sum_{s,a,h}\mathbb{P}\left(\exists n\in\{1,2,\ldots,K\}:\sum_{s^{\prime}}\left|P_{h}^{*}(s^{\prime}|s,a)-\frac{\sum_{i=1}^{n}\mathbbm{1}\{X_{i}(s,a)=s^{\prime}\}}{n}\right|>\frac{2b_{h}^{n\vee\lceil K^{\gamma}\rceil}(n)}{H-h-1}\right)$
$\leq\sum_{s,a,h}\sum_{n=1}^{K}\mathbb{P}\left(\sum_{s^{\prime}}\left|P_{h}^{*}(s^{\prime}|s,a)-\frac{\sum_{i=1}^{n}\mathbbm{1}\{X_{i}(s,a)=s^{\prime}\}}{n}\right|>\frac{2b_{h}^{n\vee\lceil K^{\gamma}\rceil}(n)}{H-h-1}\right)$ (union bound)
$\leq\sum_{s,a,h}\sum_{n=1}^{K}2^{S}\exp\left(-\frac{n}{2}\left(\frac{2b_{h}^{n\vee\lceil K^{\gamma}\rceil}(n)}{H-h-1}\right)^{2}\right)$ ((Weissman et al., 2003, Theorem 2.1))
$\leq SAHK\begin{cases}\displaystyle\ \exp\left(-2\mu K^{\alpha}\right),&\text{{KD}},\\[8.61108pt] \displaystyle\ \exp\left(-2\mu K^{\gamma\alpha}\right),&\text{{KI}}.\end{cases}\eqqcolon\delta_{K}^{\mathrm{mb}}(\gamma)$ (Lemma 1)
where in step (a), $X_{i}(s,a)\overset{\text{i.i.d.}}{\sim}P_{h}^{*}(\cdot\mid s,a)$, and the inequality follows from the facts that $n_{h}^{k}(s,a)\leq K$, a coupling argument, and that for any $n\leq k$ with $k\geq K^{\gamma}$ we have $b_{h}^{n\vee\lceil K^{\gamma}\rceil}(n)<b_{h}^{k}(n)$.
Decomposing the tail probability of the regret into the contribution of the bad event and its complement, we have
$\mathbb{P}(R_{K}\geq x)$ $\leq\mathbb{P}(R_{K}^{(1)}\geq x-\lceil K^{\gamma}\rceil H)$
$\leq\mathbb{P}\left(\left(\bar{\mathcal{G}}^{K}\right)^{\mathrm{c}}\right)+\mathbb{P}\left(\left\{R_{K}^{(1)}\geq x-\lceil K^{\gamma}\rceil H\right\}\cap\bar{\mathcal{G}}^{K}\right)$
$\leq\delta_{K}^{\mathrm{mb}}(\gamma)+\mathbb{P}\left(\left\{R_{K}^{(1)}\geq x-\lceil K^{\gamma}\rceil H\right\}\cap\bar{\mathcal{G}}^{K}\right)$
$=\delta_{K}^{\mathrm{mb}}(\gamma)+\mathbb{P}\left(R_{K}^{(1)}\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}\geq x-\lceil K^{\gamma}\rceil H\right).$ (since $x>\lceil K^{\gamma}\rceil H$)
We proceed with bounding the accumulative regret $R_{K}^{(1)}=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left(V_{0}^{*}(s_{0}^{k})-V_{0}^{\bm{\pi}_{k}}(s_{0}^{k})\right)$, under the event $\bar{\mathcal{G}}^{K}$.
$R_{K}^{(1)}\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$
$=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left[\sum_{h=0}^{H-1}\mathrm{clip}\left(\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})\right)-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ (by (5))
$=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left[\sum_{h=0}^{H-1}\mathrm{clip}\left(\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})\right)\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$
$=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\Bigg[\sum_{h=0}^{H-1}\mathrm{clip}\left(V_{h}^{*}(s_{h}^{k})-V_{h}^{k}(s_{h}^{k})+V_{h}^{k}(s_{h}^{k})-Q_{h}^{*}(s_{h}^{k},a_{h}^{k})\right)\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}\mathbbm{1}\{\mathcal{G}^{k}\}$
$\qquad\qquad\qquad\qquad-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\Bigg]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$
$\leq\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\Bigg[\sum_{h=0}^{H-1}\mathrm{clip}\left(V_{h}^{*}(s_{h}^{k})-V_{h}^{k}(s_{h}^{k})+V_{h}^{k}(s_{h}^{k})-V_{h}^{\bm{\pi}^{k}}(s_{h}^{k})\right)\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}\mathbbm{1}\{\mathcal{G}^{k}\}$
$\qquad\qquad\qquad\qquad-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\Bigg]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ ($V_{h}^{\bm{\pi}^{k}}(s_{h}^{k})=Q_{h}^{\bm{\pi}^{k}}(s_{h}^{k},a_{h}^{k})\leq Q_{h}^{*}(s_{h}^{k},a_{h}^{k})$)
$\leq\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left[\sum_{h=0}^{H-1}\mathrm{clip}\left(V_{h}^{k}(s_{h}^{k})-V_{h}^{\bm{\pi}^{k}}(s_{h}^{k})\right)\mathbbm{1}\{\mathcal{G}^{k}\}\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ (by (6))
$=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left[\sum_{h=0}^{H-1}\mathrm{clip}\left(\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}\right)\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\},$ (7)
where $\tilde{\Delta}_{h}^{k}(s_{h}^{k})\coloneqq V_{h}^{k}(s_{h}^{k})-V_{h}^{\bm{\pi}^{k}}(s_{h}^{k})$. We express $\tilde{\Delta}_{h}^{k}(s_{h}^{k})$ recursively, introducing terms that can be bounded either via concentration inequalities for $\hat{\bm{P}}_{k}$ around $\bm{P}^{*}$ or by the bonus functions.
$\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}$ $\leq\mathbbm{1}\{\mathcal{G}^{k}\}b_{h}^{k}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))+\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{s^{\prime}}\left(\hat{P}^{k}_{h}(s^{\prime}\mid s_{h}^{k},a_{h}^{k})-P^{*}_{h}(s^{\prime}\mid s_{h}^{k},a_{h}^{k})\right)V_{h+1}^{k}(s^{\prime})$
$\qquad+\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{s^{\prime}}P^{*}_{h}(s^{\prime}\mid s_{h}^{k},a_{h}^{k})\left(V_{h+1}^{k}(s^{\prime})-V_{h+1}^{\bm{\pi}^{k}}(s^{\prime})\right)$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\mathbbm{1}\{\mathcal{G}^{k}\}b_{h}^{k}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))+\frac{H-h-1}{2}\mathbbm{1}\{\mathcal{G}^{k}\}\|P_{h}^{*}(\cdot\mid s_{h}^{k},a_{h}^{k})-\hat{P}_{h}^{k}(\cdot\mid s_{h}^{k},a_{h}^{k})\|_{1}$
$\qquad+\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{s^{\prime}}P^{*}_{h}(s^{\prime}\mid s_{h}^{k},a_{h}^{k})\tilde{\Delta}_{h+1}^{k}(s^{\prime})$
$\leq 2\mathbbm{1}\{\mathcal{G}^{k}\}b^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))+\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{s^{\prime}}P^{*}_{h}(s^{\prime}\mid s_{h}^{k},a_{h}^{k})\tilde{\Delta}_{h+1}^{k}(s^{\prime})$
$=2\mathbbm{1}\{\mathcal{G}^{k}\}b^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))+\mathbb{E}\left[\mathbbm{1}\{\mathcal{G}^{k}\}\tilde{\Delta}_{h+1}^{k}(s_{h+1}^{k})\mid\mathcal{F}_{h}^{k}\right]$
$=2\mathbbm{1}\{\mathcal{G}^{k}\}b^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))+\tilde{\Delta}_{h+1}^{k}(s_{h+1}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}-\tilde{\eta}_{h+1}^{k}(s_{h+1}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}.$ (8)
where $(a)$ follows from Hölder’s inequality, and noting that $\sum_{s}(P_{h}^{*}(s\mid s_{h}^{k},a_{h}^{k})-\hat{P}_{h}^{k}(s\mid s_{h}^{k},a_{h}^{k}))=0$, and the sequence $\{\tilde{\eta}^{k}_{h}(s_{h}^{k})\}_{h=1,k=0}^{H-1,K-1}$ defined as
$\tilde{\eta}^{k}_{h}(s_{h}^{k})=\tilde{\Delta}^{k}_{h}(s_{h}^{k})-\mathbb{E}\left[\tilde{\Delta}^{k}_{h}(s_{h}^{k})\big|\mathcal{F}_{h-1}^{k}\right],$
is a Martingale difference sequence with the filtration $\{\mathcal{F}_{h-1}^{k}\}_{h=1,k=0}^{H-1,K-1}$. For the sake of notational simplicity, we define $\tilde{\eta}^{k}_{H}\equiv 0$ for all $k\in[K]$. Summing (8) over $\{h,h+1,\cdots,H-1\}$, the terms telescope, yielding
$\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}\leq 2\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))-\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}\tilde{\eta}_{l+1}^{k}(s_{l+1}^{k}).$ (9)
Next, the claim the following inequality holds:
$\mathrm{clip}\left(\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}\right)\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}$
$\qquad\leq\,\mathrm{clip}\left(4\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))\right)-2\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}\tilde{\eta}_{l+1}^{k}(s_{l+1}^{k}).$ (10)
If $a_{h}^{k}\in\mathcal{A}_{h}^{*}(s_{h}^{k})$, the inequality holds trivially. Suppose that $a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})$. Note that,
$\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}$ $=(V^{k}_{h}(s_{h}^{k})-V_{h}^{\bm{\pi}^{k}}(s_{h}^{k}))\mathbbm{1}\{\mathcal{G}^{k}\}$
$\geq(V^{*}_{h}(s_{h}^{k})-Q_{h}^{*}(s_{h}^{k},a_{h}^{k}))\mathbbm{1}\{\mathcal{G}^{k}\}$ (by $\mathbbm{1}\{\mathcal{G}^{k}\}$ and $V_{h}^{\bm{\pi}^{k}}(s_{h}^{k})\leq Q_{h}^{*}(s_{h}^{k},a_{h}^{k})$)
$=\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}$
$\geq\mathrm{gap}^{*}\mathbbm{1}\{\mathcal{G}^{k}\}$ ($a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})$)
which implies $\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}=\mathrm{clip}\left(\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}\right)$ given $a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})$. Now, we have the following cases:
  1. 1.
    If $\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}\tilde{\eta}_{l+1}^{k}(s_{l+1}^{k})\geq 2\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))$, then
    $\mathrm{clip}\left(\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}\right)$ $=\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}$
    $\leq-2\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}\tilde{\eta}_{l+1}^{k}(s_{l+1}^{k})$ (by (9))
  2. 2.
    If $2\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))>-\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}\tilde{\eta}_{l+1}^{k}(s_{l+1}^{k})$, then
    $\mathrm{gap}^{*}\mathbbm{1}\{\mathcal{G}^{k}\}\leq\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}$ $\leq 4\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k})),$ (by (9))
    which implies
    $\mathrm{clip}\left(4\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))\right)=4\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k})).$
Substituting (10) into (7) yields
$R_{K}^{(1)}\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ $\leq\,\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\Bigg[\sum_{h=0}^{H-1}\mathrm{clip}\left(4\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))\right)$
$\qquad\qquad-2\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{h=0}^{H-1}\sum_{l=h}^{H-1}\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}\tilde{\eta}_{l+1}^{k}(s_{l+1}^{k})-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\Bigg]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$
$\leq\,\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\Bigg[\sum_{h=0}^{H-1}\sum_{l=h}^{H-1}\mathrm{clip}\left(4Hb^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))\right)$ (by (4))
$\qquad\qquad-2\sum_{h=0}^{H-1}\sum_{l=h}^{H-1}\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}\tilde{\eta}_{l+1}^{k}(s_{l+1}^{k})-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\Bigg]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$
$=\,\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\Bigg[\sum_{h=0}^{H-1}(h+1)\mathrm{clip}\left(4Hb^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))\right)$
$\qquad\qquad-\sum_{h=0}^{H-1}\left(2\left(\sum_{l=0}^{h}\mathbbm{1}\{a_{l}^{k}\notin\mathcal{A}_{l}^{*}(s_{l}^{k})\}\right)\tilde{\eta}_{h+1}^{k}(s_{h+1}^{k})+\eta^{k}_{h+1}(s_{h+1}^{k})\right)\Bigg]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$
$=\,\left[\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}(h+1)\mathrm{clip}\left(4Hb^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))\right)-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\bar{\eta}_{h+1}^{k}\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ (11)
where the sequence $\{\bar{\eta}_{h}^{k}\}_{h=1,k=0}^{H-1,K-1}$ defined as
$\bar{\eta}^{k}_{h+1}=2\left(\sum_{l=0}^{h}\mathbbm{1}\{a_{l}^{k}\notin\mathcal{A}_{l}^{*}(s_{l}^{k})\}\right)\tilde{\eta}_{h+1}^{k}(s_{h+1}^{k})+\eta^{k}_{h+1}(s_{h+1}^{k}),$
is a Martingale difference sequence with the filtration $\{\mathcal{F}_{h-1}^{k}\}_{h=1,k=0}^{H-1,K-1}$, and satisfies $|\bar{\eta}^{k}_{h+1}|\leq 5H^{2}$.
Next, we bound the first term in (11). We have
$\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}(h+1)\mathrm{clip}\left(4Hb^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))\right)$ $\leq H\sum_{k=0}^{K-1}\sum_{h=0}^{H-1}\mathrm{clip}\left(4Hb^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))\right)$
$\leq H\sum_{k=0}^{K-1}\sum_{h=0}^{H-1}\mathrm{clip}\left(4H^{2}\sqrt{\frac{0.5S\ln 2+\mu K^{\alpha}}{\max\{1,n_{h}^{k}(s_{h}^{k},a_{h}^{k})\}}}\right)$
$\leq H\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\sum_{h=0}^{H-1}\sum_{n=1}^{\infty}\mathrm{clip}\left(4H^{2}\sqrt{\frac{0.5S\ln 2+\mu K^{\alpha}}{\max\{1,n\}}}\right)$
$=H\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\sum_{h=0}^{H-1}\sum_{n=1}^{\bar{n}}\left(4H^{2}\sqrt{\frac{0.5S\ln 2+\mu K^{\alpha}}{\max\{1,n\}}}\right)$
$\leq 8H^{4}SA\sqrt{0.5S\ln 2+\mu K^{\alpha}}\sqrt{\bar{n}}$
$\leq\frac{32H^{6}SA(0.5S\ln 2+\mu K^{\alpha})}{\mathrm{gap}^{*}}\eqqcolon m_{K}^{\mathrm{mb}}$ (12)
where
$\bar{n}=\left\lfloor\frac{16H^{4}(0.5S\ln 2+\mu K^{\alpha})}{{\mathrm{gap}^{*}}^{2}}\right\rfloor.$
Combining the bounds, we have
$\mathbb{P}\left(\left\{R_{K}^{(1)}\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}\geq x-\lceil K^{\gamma}\rceil H\right\}\right)$
$\qquad\qquad\leq\mathbb{P}\left(\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}\left(m_{K}^{\mathrm{mb}}-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\bar{\eta}_{h+1}^{k}\right)\geq x-\lceil K^{\gamma}\rceil H\right)$
$\qquad\qquad\leq\mathbb{P}\left(-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\bar{\eta}_{h+1}^{k}\geq x-\lceil K^{\gamma}\rceil H-m_{K}^{\mathrm{mb}}\right)$ (since $x>\lceil K^{\gamma}\rceil H$)
$\qquad\qquad\leq\exp\left(-\frac{(x-\lceil K^{\gamma}\rceil H-m_{K}^{\mathrm{mb}})^{2}}{50H^{5}K}\right).$
where the last inequality follows by applying Azuma–Hoeffding.
Proof of Proposition 1.
We have
$\mathbb{E}[R_{K}]=$ $\mathbb{E}[R_{K}^{(0)}]+\mathbb{E}[R_{K}^{(1)}]$
$\leq$ $H\lceil K^{\gamma}\rceil+\mathbb{E}\left[R_{K}^{(1)}\left(\mathbbm{1}\left\{\bar{\mathcal{G}}^{K}\right\}+\mathbbm{1}\left\{\left(\bar{\mathcal{G}}^{K}\right)^{\mathrm{c}}\right\}\right)\right]$
$=$ $H\lceil K^{\gamma}\rceil+\mathbb{E}\left[R_{K}^{(1)}\mathbbm{1}\left\{\bar{\mathcal{G}}^{K}\right\}\right]+\mathbb{E}\left[R_{K}^{(1)}\mathbbm{1}\left\{\left(\bar{\mathcal{G}}^{K}\right)^{c}\right\}\right]$
$\leq$ $H\lceil K^{\gamma}\rceil+\mathbb{E}\left[R_{K}^{(1)}\mathbbm{1}\left\{\bar{\mathcal{G}}^{K}\right\}\right]+H(K-\lceil K^{\gamma}\rceil)\mathbb{P}\left(\left(\bar{\mathcal{G}}^{K}\right)^{\mathrm{c}}\right).$ (13)
Note that by (11) and (12), we have
$R_{K}^{(1)}\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}\leq\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}\left(m_{K}^{\mathrm{mb}}-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\bar{\eta}_{h+1}^{k}\right)$
Hence, we obtain
$\mathbb{E}\left[R_{K}^{(1)}\mathbbm{1}\left\{\bar{\mathcal{G}}^{K}\right\}\right]$ $\leq m_{K}^{\mathrm{mb}}+\mathbb{E}\left[\left(-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\bar{\eta}_{h+1}^{k}\right)\mathbbm{1}\left\{\bar{\mathcal{G}}^{K}\right\}\right]$
$=m_{K}^{\mathrm{mb}}+\mathbb{E}\left[-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\bar{\eta}_{h+1}^{k}\right]+\mathbb{E}\left[\left(\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\bar{\eta}_{h+1}^{k}\right)\mathbbm{1}\left\{\left(\bar{\mathcal{G}}^{K}\right)^{\mathrm{c}}\right\}\right]$
$\leq m_{K}^{\mathrm{mb}}+5KH^{3}\delta_{K}^{\mathrm{mb}}(\gamma).$ (14)
Combining (13) and (14), we get
$\mathbb{E}[R_{K}]\leq H\lceil K^{\gamma}\rceil+m_{K}^{\mathrm{mb}}+6KH^{3}\delta_{K}^{\mathrm{mb}}(\gamma).$
Lemma 1.
The following bound holds:
$\sum_{n=1}^{K}$ $\exp\left(-2n\left(\frac{b_{h}^{n\vee\lceil K^{\gamma}\rceil}(n)}{H-h-1}\right)^{2}\right)\leq 2^{-S}\times\begin{cases}\displaystyle K\exp\left(-2\mu{K^{\alpha}}\right),&\text{{KD}},\\[8.61108pt] \displaystyle K\exp\left(-2\mu K^{\gamma\alpha}\right),&\text{{KI}}.\end{cases}$
Proof of Lemma 1.
Recall that the function $b^{k}_{h}(\cdot)$ is defined as
$b^{k}_{h}(n)=\begin{cases}(H-h-1)\sqrt{\dfrac{0.5S\ln 2+\mu K^{\alpha}}{n}},&\text{{KD}},\\[8.61108pt] (H-h-1)\sqrt{\dfrac{0.5S\ln 2+\mu(k+1)^{\alpha}}{n}},&\text{{KI}},\end{cases}\qquad b^{k}_{h}(0)\coloneqq\infty.$
Let $S_{K}$ denote the summation in Lemma 1, i.e.,
$S_{K}\coloneqq\sum_{n=1}^{K}\exp\left(-2n\left(\frac{b_{h}^{n\vee\lceil K^{\gamma}\rceil}(n)}{H-h-1}\right)^{2}\right).$
The analysis depends on the choice of $b^{k}_{h}(\cdot)$.
  • KD: We have
    $S_{K}$ $\leq\sum_{n=1}^{K}\exp\left(-S\ln 2-2\mu K^{\alpha}\right)=2^{-S}K\exp\left({-2\mu K^{\alpha}}\right)$
  • KI: Define $S^{(1)}_{K}$ and $S^{(2)}_{K}$ as follows:
    $S^{(1)}_{K}\coloneqq\sum_{n=1}^{\lceil K^{\gamma}\rceil}\exp\left(-2n\left(\frac{b_{h}^{\lceil K^{\gamma}\rceil}(n)}{H-h-1}\right)^{2}\right),\qquad S^{(2)}_{K}\coloneqq\sum_{n=\lceil K^{\gamma}\rceil+1}^{K}\exp\left(-2n\left(\frac{b_{h}^{n}(n)}{H-h-1}\right)^{2}\right).$
    Notice that $S_{K}=S^{(1)}_{K}+S^{(2)}_{K}$, and that
    $b_{h}^{\lceil K^{\gamma}\rceil}(n)=(H-h-1)\sqrt{\frac{0.5S\ln 2+\mu(\lceil K^{\gamma}\rceil+1)^{\alpha}}{n}},\quad b_{h}^{n}(n)=(H-h-1)\sqrt{\frac{0.5S\ln 2+\mu(n+1)^{\alpha}}{n}}.$
    We proceed by bounding $S_{K}^{(1)}$ and $S_{K}^{(2)}$. We begin with $S_{K}^{(1)}$, for which we have
    $S_{K}^{(1)}$ $=\sum_{n=1}^{\lceil K^{\gamma}\rceil}\exp\left(-2n\left(\frac{0.5S\ln 2+\mu(\lceil K^{\gamma}\rceil+1)^{\alpha}}{n}\right)\right)\leq 2^{-S}\lceil K^{\gamma}\rceil\exp\left(-2\mu K^{\gamma\alpha}\right)$
We next bound $S_{K}^{(2)}$:
$S^{(2)}_{K}$ $=\sum_{n=\lceil K^{\gamma}\rceil+1}^{K}\exp\left(-2n\left(\frac{b_{h}^{n}(n)}{H-h-1}\right)^{2}\right)$
$\leq\sum_{n=\lceil K^{\gamma}\rceil+1}^{K}2^{-S}\exp\left(-2\mu n^{\alpha}\right)$
$\leq 2^{-S}(K-\lceil K^{\gamma}\rceil)\exp\left(-2\mu K^{\gamma\alpha}\right).$
Putting everything together, we obtain the desired result. ∎